User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

AuScope Operations Meeting Thursday 21st April, held at 12pm.

Chair: Ellen

Apologies: Nick

Attendees: Lucia, Jim, Jesse, Bryn, Ross, Jon, Tiege, Lucas, Arwin, Liz

Current Issues

  • Problems with drive hydraulics on 26m, cancelled today's experiment (RD1603).
    1. New something being bought by Eric
  • New PCFS at Hobart is working great (thanks Jamie). Has anyone had any problems?
  • Maser maintenance at both Katherine and Yarragadee currently, hence Brett is not at the Observatory. Warren is at Mt Pleasant Tuesdays to Thursdays and occasionally Mondays.
  • CRDS82 - “still getting ve error: ERROR ch -308 ve total power integrator overflow” - Bryn, what was this and was it fixed?
    1. High tsys values due to RFI
  • R4734 - Yarragadee: consistent offset between mark 5 and other computers of 0.1 - 0.4 visible in fmset with no sync error, could not be fixed. Caused by power failure affecting GPS receiver (new UPS installed) - had to change to Hobart's. Fixed itself 4+ hours later. Is there any way to fix this other than hope?
  1. Probably ntp problem
  • R1735 - A couple of instances where recorder started ~ 7-10 seconds before onsource at both Katherine and Yarragadee.
    1. possibly due to tag along experiment
  • High system temperatures - run systemp12 again to make sure it wasn't just one dodgy reading. Monitor and make note of how many scans were affected by consistent high tsys values.
    1. elevation and weather affects tsys values
    2. would be cool if monica can track changes
  • Anything else to note?
  • VSNs all reported as the same for 4 telescopes in end message (Jesse check experiment) - keep an eye on this
/home/www/auscope/opswiki/data/pages/operations/agenda2016_04_21.txt · Last modified: 2016/04/21 02:17 by Ellen Manning